# Humanoid Robots and Humanoid AI: Review, Perspectives and Directions

### v2

@@ 1. Introduction
Despite these promising projections and fast-paced developments, only a limited number of AI humanoids are empowered by LLMs or GenAI, and none actually hold human-level AI and human intelligence like humanity. This discrepancy is attributed not only to the so-called “uncanny valley effect” (Mori et al., 2012) but more fundamentally to the “humanoid humanity dilemma” (or more broadly, machine humanity dilemma). Humanoid humanity dilemma refers to the plight of empowering human-looking humanoids with realistic to natural humanness. Both the uncanny valley effect and the humanoid humanity dilemma highlight significant untapped potential and significant gaps within the realm of humanoid AI, particularly in the foundation and design of realistic humanoids. Accordingly, discussion and speculation about the future trajectory of humanoid AI have to answer a foundational question: What does the evolution and future hold for AI humanoids? This has been intensified over the rapid evolution from ANI to AGI or even artificial superintelligence (ASI) (Yampolskiy, 2015) as well as other specialties including metaverse and decentralized AI (Cao, 2022b; Li et al., 2024).
@@ 2.2. Humanization Paradigm Shift: Human-looking, Human-like and Human-level Humanoids
First, human-looking humanoids are mounted with humanoid structures to make humanoid look like human beings, with some having still or moving attributes. As shown in Table LABEL:tab:humanoids, most of humanoids are equipped with simple and abstract human-looking parts, while robots like Geminoid and Junko Chihira are with more nuanced facial and head features. Humanoids may be driven by wheel and bipedal facilities. Then, human-like humanoids further humanize robots with representative human traits and functions. Humanoid senses, behaviors and functions make human-like humanoids not only look humanly but also behave and function humanly. Such humanoids tend to support human-friendly immersive social interactions, such as by nuanced eye contact, and expressive facial and and postural response. Lastly, human-level humanoids narrow the gaps between humanoids being machine with machine intelligence and the clone (‘doppelganger’) of humans with some of simulated human uniqueness and identity of humanity and intelligence.
@@
There are about 30 types of humanoid robots available in the market and literature, while this number is rapidly updated. Table LABEL:tab:humanoids in Appendix provides an overview and comparison of 30 humanoid robots identifiable in public resources, which are endowed with certain human-like features and technical and AI-enabled functions. From the AI capacity perspective, we categorize existing humanoids into 1) naive human-looking humanoids with some human appearance but limited to none AI functions, 2) ANI-driven standalone humanoids with limited ANI functions on board only, and 3) GenAI-enabled networked humanoids with on-humanoid and cloud-based operations and more capable AI functions in particular GenAI capabilities like Ameca with real-time connection to ChatGPT. Humanoids at these three capacity phases do not reach AGI capabilities on humanoid or on cloud.
@@ 3.1. Functional Specifications of Humanoids
Functional specifications of human-like and lifelike humanoids define and enable how an AI humanoid look like, what it does, and how it implements its purposes, tasks, and operations, etc. They determine what humanoid structures, senses, behaviors and functions to implement and operate, as summarized in Table LABEL:tab:functionalspecifications. While humanoids differ in their functions and none holds all functions, humanoids are expected to satisfy functional specifications defining mechanical, electrical, biological and social designs with anthropomorphic and mimetic features. Such humanoids imitate human intelligence by their enabling appearance, brain, perceptiveness, activeness, connectiveness, interaction, organization, knowledge, cybernetics, generalization, and system.
@@ 3.1.4. Humanoid Activeness: Behaviors, Motions and Dynamics
Humanoid activeness are embodied through their behaviors, kinematics, and dynamics. The dynamics of human systems, behaviors, and intelligence are fundamental for human functions, operations, and evolution. Understanding human behaviors and dynamics can inform and transform humanoid behaviors and dynamics. Humanoid behaviors capture how a humanoid behaves, acts, and reacts, etc., supporting various forms of operations, such as gesture, movement, and interactions with other robots or humans (Bärmann et al., 2023). Behaviors emerge over the lifecycle of humanoid tasking, generating sequential and full-cycle state spaces with behavior attributes, processes, consequences, and impacts (Cao, 2010). For example, humanoid gestures (Carfì and Mastrogiovanni, 2023) comprise human gesture-simulated appearances, behaviors, and states, enabling humanoid’s facial expression, eye gaze, communicative pointing and signs, and manipulative motions, etc. Humanoid movements embody actions such as eye, hand, and leg movement. These low-level behaviors further enable high-level humanoid motions, implementing general behavioral functions and purposes such as communication, engagement, grasping, navigation, and teleoperation. Under specific tasks and scenarios, humanoids may take imitative, tactile, social, walking, grasping, acrobatic, or amphibious behaviors, which may be stylistic, personalized, safe or hazardous.
@@ 3.1.5. Humanoid Connectiveness: Networking and Communication
LoC-driven humanoid communications expand linguistic communications from speech, message-passing, and dialogue to query-answering, retrieval, and recommendations; and non-linguistic communications from vision (Juang, 2024) and camera/eye contact to responsive emotions, and postural, gestural, behavioral, or graphical interactions or movements. These transform unimodal to multimodal LoC with language-based to multimodal communications. Non-expressive communications such as through sensory information consumption and sharing are complemented by expressive communications such as emotional, sentimental, affective, or behavioral expressions and exchange of facial expressions, spoken emotions, textual sentiments, or behaviors such as gestures, postures, body language, and actions. Besides multimodalities, multimodal LoC also incorporates multimodal transformation and translation in bilateral manners, such as via text-to-speech translation, speech-to-text translation, image-to-text transformation, text-to-image transformation, vision-to-language transformation, and language-to-vision transformation; multilateral, such as vision-language-to-action (VLA) transformation, or vision-language-to-emotion (VLE) transformation (Li et al., 2024); or mechanism-based, e.g., demonstration-to-action, or imitation-to-action. These translation and transformation tasks undertake cross-modality or cross-mechanism information transport, which may take place in an invasive, non-invasive, or hybrid mode.
@@ 3.1.7. Humanoid Organization: Planning and Search
With humanoids enabled by learning advances including LLMs and LMMs, learn-to-plan, or planning by learning (Ganesan et al., 2024), becomes a new fashion of humanoid planning in complex contexts. Humanoids navigate, search, respond and act per justifications, predictions, recommendations, or retrieved findings by techniques including probabilistic roadmap, evolutionary computing, geometric modeling, reinforcement learning, and deep learning. Learn-to-plan is particularly important for interacting with, manipulating, or responding to unseen objects, scenarios, tasks in semi-open to open worlds. For example, a robot enabled by a pretrained VLM can extract object category-specific features from images and texts (Stone et al., 2023). Further, a robot can undertake actions or complete instructions with policies informed or constrained by categorized objects with imagery and textual instructions, such as SayCan (Ichter et al., 2022) which plans-to-act based on input language. Google RT-2 (Brohan et al., 2023b) is another VLM-based learn-to-plan advance, which builds on a VLM backbone, plans from both image and text commands, and conducts visually grounded planning. Its chain-of-thought reasoning can enable learning long-horizon planning and low-level skills to achieve highly-improved robotic policies by instantiating VLAs based on PaLM-E (Driess et al., 2023) and PaLI-X (Chen et al., 2023). RT-2 can significantly improve generalization performance and emergent capabilities using the power of web-scale vision-language pretraining (VLP) (Agrawal et al., 2022).
@@ 3.1.8. Humanoid Processing: Representation, Analytics, Learning, Reasoning and Recommendation
Humanoids are associated with their data, behaviors and environments. Humanoids represent, analyze, learn, discover, reason about and recommend them to knowledge to empower humanoid functions, operations, and performance. Humanoid knowledge enables their capabilities and quality of services through analytics and learning from data, behaviors, and environments of themselves or other humanoids, robots, humans, and objects. They undertake goals and tasks to gain and strengthen their specific to comprehensive humanoid intelligence, particularly by conducting machine vision, reasoning, reinforcement learning, and learning by demonstrations, imitations, interactions, behaviors, and feedback. As illustrated on learning to communicate and interact in Sections 3.1.5 and 3.1.6, autonomous and intelligent humanoids perform on-device and decentralized to cloud and online, shallow to deep, small to scalable, pretrained to finetuned, unsupervised to supervised, unimodal to multimodal, uni-task to multi-task, on-policy to off-policy, and specific to general learning in closed to open settings or worlds. In particular, generative AI especially LLMs and LMMs enables more scalable, self-supervised, scenario-based, and real-time humanoid intelligence, such as for real-time, interactive, affective multiparty conversation (Li et al., 2024), which was impossible to achieve before.
@@
Another humanoid learning paradigm is transformative/translational humanoid learning (THL), where humanoids learn to transform or translate. THL progresses from vision-language modeling (VLM) (Stone et al., 2023; Kawaharazuka et al., 2024) to vision-language-to-action (VLA) (Brohan et al., 2023b), vision-language-to-emotion (VLE) (Li et al., 2024), and vision-language-to-behavior (VLB) modeling. These foster new directions of multimodal modeling to actions (MMA) and multimodal modeling to decisions (MMD) for actionable humanoid learning. In Section 4, we will further discuss humanoid analytics and directions of incorporating LLMs, GenAI and deep learning into humanoids, and specifically on opportunities including omnimodal perception-to-behavior and mind-to-action modeling for mindful, cognitive, and actionable humanoids in Section 4.4.
@@ 3.2. Nonfunctional Specifications of Humanoids
With the fast growth and advancement of robotic techniques and functions, humanoid robots are expected safe and secure in their operations and action-taking; trustful and responsible for their actions; transparent and explainable in their outputs; empathetic and rational during teaming, collaboration and interaction; and compliant, legal and ethical for integrity and regulation. Accordingly, the functions and quality of humanoids need to be ensured by various nonfunctional requirements, objectives and constraints. As summarized in Table LABEL:tab:nonfunctionalspecifications, they cover many aspects from traditional requirements on the robotic quality of services to new control problems relating to humanized requirements. We categorize such nonfunctional specifications into four aspects below: 1) human likeness satisfaction, 2) humanoid capability maturity, 3) humanoid performance measurement, and 4) humanoid impact estimation.
@@ 3.2.2. Humanoid Capability Maturity
Humanoid functional proficiency refers to the degree of competence and skills of a humanoid in performing their tasks and functions illustrated in Section 3.1. Humanoid functional proficiency can be customized into cognitive proficiency of implementing humanoid brain, physical proficiency of implementing anthropomorphism, perceptive proficiency of performing vision, speech, hearing and other sensory functions, behavioral proficiency of achieving activeness, connective proficiency of performing networking and communications, interactive proficiency of conducting collaboration and competition, organizational proficiency of planning and search, knowledge proficiency of representation, analytics, learning, reasoning and recommendation, cybernetic proficiency of conducting manipulation and control, generalizable proficiency of achieving adaptation and optimization, and system proficiency of humanoid design and implementation. Humanoid proficiency may also be categorized into task-specific proficiency in undertaking a specific task, nonfunctional proficiency and other specific types of proficiency (Norton et al., 2022), such as emotional proficiency of demonstrating empathy, and social proficiency of engaging humans and adapting to social, ethical and cultural norms. Most humanoids as shown in Table LABEL:tab:humanoids only possess basic to intermediate levels of proficiency of some of these aspects. For example, few humanoids like Ameca have the potential of incorporating cognitive, sentimental or emotional states and features, such as facial expressions. In addition, social humanoids may hold humanized features and traits, such as personality as embodied by extroversion, agreeableness, conscientiousness, neuroticism and openness (Fong et al., 2003).
@@ 4. Realistic Humanoids: Other Perspectives
The comprehensive analysis of existing humanoids in Section 2, as outlined in Table LABEL:tab:humanoids, and the systematic discussion on functional and nonfunctional specifications of humanoids in Section 3 disclose various gaps, limitations of existing work and the implications for building new-generation humane humanoids. Fig. 2 highlights other perspectives and directions in developing humanoid AI: humanoid humanity, mind-to-action, biomedicine, animation, virtual-real humanoid mapping and symbiosis, metaverse, humanoid-human symbiosis, generation, decentralization, standardization, and quality of services. With space limitation, this section expands Section 3 to several other perspectives and directions: cultivating humanoid humanity and subjectivity, building mind-to-action mindful and actionable humanoids, supporting transitional omnimodal humanoid modeling, and hybridizing virtual-real humanoid-human-metaverse symbiosis.
@@ 4.4. Transitional Omnimodal Humanoid Modeling
In recent years, an increasing attention has been paid to multimodal deep learning and specifically LMMs, such as visual query-answering (VQA), VLMs, VLP, image-language pretraining, image-to-text generation, and text-to-image generation. Building on Google Robotic Transformer 1 (RT-1) on multitask modeling (Brohan et al., 2023a), Google Robotic Transformer 2 (RT-2) (Brohan et al., 2023b) further supports high-capacity VLA modeling on web data and robotic vision data, and translates the learned knowledge into generalized instructions to control robotic actions. RT-2 uses the pathways language and image model (PaLI-X) and pathways language model embodied (PaLM-E) as the backbone.

### v3

@@ 1. Introduction
Despite these promising projections and fast-paced developments, only a limited number of AI humanoids are empowered by LLMs or GenAI, and none actually hold human-level AI and human intelligence like humanity. This discrepancy is attributed not only to the so-called “uncanny valley effect” (Mori et al., 2012) but more fundamentally to the “humanoid humanity dilemma” (or more broadly, machine humanity dilemma). Humanoid humanity dilemma refers to the plight of empowering human-looking humanoids with realistic to natural humanness. Both the uncanny valley effect and the humanoid humanity dilemma highlight significant untapped potential and significant gaps within the realm of humanoid AI, particularly in the foundation and design of realistic humanoids. Accordingly, discussion and speculation about the future trajectory of humanoid AI have to answer a foundational question: What does the evolution and future hold for AI humanoids? This has been intensified over the rapid evolution from ANI to AGI or even artificial superintelligence (ASI) (Yampolskiy, 2015) as well as other specialties including metaverse and decentralized AI (Cao, 2022b; Li et al., 2025).
@@ 2.2. Humanization Paradigm Shift: Human-looking, Human-like and Human-level Humanoids
First, human-looking humanoids are mounted with humanoid structures to make humanoid look like human beings, with some having still or moving attributes. As shown in Table 2 (Cao, 2025), most of humanoids are equipped with simple and abstract human-looking parts, while robots like Geminoid and Junko Chihira are with more nuanced facial and head features. Humanoids may be driven by wheel and bipedal facilities. Then, human-like humanoids further humanize robots with representative human traits and functions. Humanoid senses, behaviors and functions make human-like humanoids not only look humanly but also behave and function humanly. Such humanoids tend to support human-friendly immersive social interactions, such as by nuanced eye contact, and expressive facial and and postural response. Lastly, human-level humanoids narrow the gaps between humanoids being machine with machine intelligence and the clone (‘doppelganger’) of humans with some of simulated human uniqueness and identity of humanity and intelligence.
@@
There are about 30 types of humanoid robots available in the market and literature, while this number is rapidly updated. Table 2 in (Cao, 2025) provides an overview and comparison of 30 humanoid robots identifiable in public resources, which are endowed with certain human-like features and technical and AI-enabled functions. From the AI capacity perspective, we categorize existing humanoids into 1) naive human-looking humanoids with some human appearance but limited to none AI functions, 2) ANI-driven standalone humanoids with limited ANI functions on board only, and 3) GenAI-enabled networked humanoids with on-humanoid and cloud-based operations and more capable AI functions in particular GenAI capabilities like Ameca with real-time connection to ChatGPT. Humanoids at these three capacity phases do not reach AGI capabilities on humanoid or on cloud.
@@ 3.1. Functional Specifications of Humanoids
Functional specifications of human-like and lifelike humanoids define and enable how an AI humanoid look like, what it does, and how it implements its purposes, tasks, and operations, etc. They determine what humanoid structures, senses, behaviors and functions to implement and operate, as summarized in Table 3 (Cao, 2025). While humanoids differ in their functions and none holds all functions, humanoids are expected to satisfy functional specifications defining mechanical, electrical, biological and social designs with anthropomorphic and mimetic features. Such humanoids imitate human intelligence by their enabling appearance, brain, perceptiveness, activeness, connectiveness, interaction, organization, knowledge, cybernetics, generalization, and system.
@@ 3.1.4. Humanoid Activeness: Behaviors, Motions and Dynamics
Humanoid activeness are embodied through their behaviors, kinematics, and dynamics. The dynamics of human systems, behaviors, and intelligence are fundamental for human functions, operations, and evolution. Understanding human behaviors and dynamics can inform and transform humanoid behaviors and dynamics. Humanoid behaviors capture how a humanoid behaves, acts, and reacts, etc., supporting various forms of operations, such as gesture, movement, and interactions with other robots or humans (Bärmann et al., 2024). Behaviors emerge over the lifecycle of humanoid tasking, generating sequential and full-cycle state spaces with behavior attributes, processes, consequences, and impacts (Cao, 2010). For example, humanoid gestures (Carfì and Mastrogiovanni, 2023) comprise human gesture-simulated appearances, behaviors, and states, enabling humanoid’s facial expression, eye gaze, communicative pointing and signs, and manipulative motions, etc. Humanoid movements embody actions such as eye, hand, and leg movement. These low-level behaviors further enable high-level humanoid motions, implementing general behavioral functions and purposes such as communication, engagement, grasping, navigation, and teleoperation. Under specific tasks and scenarios, humanoids may take imitative, tactile, social, walking, grasping, acrobatic, or amphibious behaviors, which may be stylistic, personalized, safe or hazardous.
@@ 3.1.5. Humanoid Connectiveness: Networking and Communication
LoC-driven humanoid communications expand linguistic communications from speech, message-passing, and dialogue to query-answering, retrieval, and recommendations; and non-linguistic communications from vision (Juang, 2024) and camera/eye contact to responsive emotions, and postural, gestural, behavioral, or graphical interactions or movements. These transform unimodal to multimodal LoC with language-based to multimodal communications. Non-expressive communications such as through sensory information consumption and sharing are complemented by expressive communications such as emotional, sentimental, affective, or behavioral expressions and exchange of facial expressions, spoken emotions, textual sentiments, or behaviors such as gestures, postures, body language, and actions. Besides multimodalities, multimodal LoC also incorporates multimodal transformation and translation in bilateral manners, such as via text-to-speech translation, speech-to-text translation, image-to-text transformation, text-to-image transformation, vision-to-language transformation, and language-to-vision transformation; multilateral, such as vision-language-to-action (VLA) transformation, or vision-language-to-emotion (VLE) transformation (Li et al., 2025); or mechanism-based, e.g., demonstration-to-action, or imitation-to-action. These translation and transformation tasks undertake cross-modality or cross-mechanism information transport, which may take place in an invasive, non-invasive, or hybrid mode.
@@ 3.1.7. Humanoid Organization: Planning and Search
With humanoids enabled by learning advances including LLMs and LMMs, learn-to-plan, or planning by learning (Ganesan et al., 2024), becomes a new fashion of humanoid planning in complex contexts. Humanoids navigate, search, respond and act per justifications, predictions, recommendations, or retrieved findings by techniques including probabilistic roadmap, evolutionary computing, geometric modeling, reinforcement learning, and deep learning. Learn-to-plan is particularly important for interacting with, manipulating, or responding to unseen objects, scenarios, tasks in semi-open to open worlds. For example, a robot enabled by a pretrained VLM can extract object category-specific features from images and texts (Stone et al., 2023). Further, a robot can undertake actions or complete instructions with policies informed or constrained by categorized objects with imagery and textual instructions, such as SayCan (Brohan et al., 2022) which plans-to-act based on input language. Google RT-2 (Zitkovich et al., 2023) is another VLM-based learn-to-plan advance, which builds on a VLM backbone, plans from both image and text commands, and conducts visually grounded planning. Its chain-of-thought reasoning can enable learning long-horizon planning and low-level skills to achieve highly-improved robotic policies by instantiating VLAs based on PaLM-E (Driess et al., 2023) and PaLI-X (Chen et al., 2023). RT-2 can significantly improve generalization performance and emergent capabilities using the power of web-scale vision-language pretraining (VLP) (Agrawal et al., 2022).
@@ 3.1.8. Humanoid Processing: Representation, Analytics, Learning, Reasoning and Recommendation
Humanoids are associated with their data, behaviors and environments. Humanoids represent, analyze, learn, discover, reason about and recommend them to knowledge to empower humanoid functions, operations, and performance. Humanoid knowledge enables their capabilities and quality of services through analytics and learning from data, behaviors, and environments of themselves or other humanoids, robots, humans, and objects. They undertake goals and tasks to gain and strengthen their specific to comprehensive humanoid intelligence, particularly by conducting machine vision, reasoning, reinforcement learning, and learning by demonstrations, imitations, interactions, behaviors, and feedback. As illustrated on learning to communicate and interact in Sections 3.1.5 and 3.1.6, autonomous and intelligent humanoids perform on-device and decentralized to cloud and online, shallow to deep, small to scalable, pretrained to finetuned, unsupervised to supervised, unimodal to multimodal, uni-task to multi-task, on-policy to off-policy, and specific to general learning in closed to open settings or worlds. In particular, generative AI especially LLMs and LMMs enables more scalable, self-supervised, scenario-based, and real-time humanoid intelligence, such as for real-time, interactive, affective multiparty conversation (Li et al., 2025), which was impossible to achieve before.
@@
Another humanoid learning paradigm is transformative/translational humanoid learning (THL), where humanoids learn to transform or translate. THL progresses from vision-language modeling (VLM) (Stone et al., 2023; Kawaharazuka et al., 2024) to vision-language-to-action (VLA) (Zitkovich et al., 2023), vision-language-to-emotion (VLE) (Li et al., 2025), and vision-language-to-behavior (VLB) modeling. These foster new directions of multimodal modeling to actions (MMA) and multimodal modeling to decisions (MMD) for actionable humanoid learning. In Section 4, we will further discuss humanoid analytics and directions of incorporating LLMs, GenAI and deep learning into humanoids, and specifically on opportunities including omnimodal perception-to-behavior and mind-to-action modeling for mindful, cognitive, and actionable humanoids in Section 4.4.
@@ 3.2. Nonfunctional Specifications of Humanoids
With the fast growth and advancement of robotic techniques and functions, humanoid robots are expected safe and secure in their operations and action-taking; trustful and responsible for their actions; transparent and explainable in their outputs; empathetic and rational during teaming, collaboration and interaction; and compliant, legal and ethical for integrity and regulation. Accordingly, the functions and quality of humanoids need to be ensured by various nonfunctional requirements, objectives and constraints. As summarized in Table 3 (Cao, 2025), they cover many aspects from traditional requirements on the robotic quality of services to new control problems relating to humanized requirements. We categorize such nonfunctional specifications into four aspects below: 1) human likeness satisfaction, 2) humanoid capability maturity, 3) humanoid performance measurement, and 4) humanoid impact estimation.
@@ 3.2.2. Humanoid Capability Maturity
Humanoid functional proficiency refers to the degree of competence and skills of a humanoid in performing their tasks and functions illustrated in Section 3.1. Humanoid functional proficiency can be customized into cognitive proficiency of implementing humanoid brain, physical proficiency of implementing anthropomorphism, perceptive proficiency of performing vision, speech, hearing and other sensory functions, behavioral proficiency of achieving activeness, connective proficiency of performing networking and communications, interactive proficiency of conducting collaboration and competition, organizational proficiency of planning and search, knowledge proficiency of representation, analytics, learning, reasoning and recommendation, cybernetic proficiency of conducting manipulation and control, generalizable proficiency of achieving adaptation and optimization, and system proficiency of humanoid design and implementation. Humanoid proficiency may also be categorized into task-specific proficiency in undertaking a specific task, nonfunctional proficiency and other specific types of proficiency (Norton et al., 2022), such as emotional proficiency of demonstrating empathy, and social proficiency of engaging humans and adapting to social, ethical and cultural norms. Most humanoids as shown in Table 2 (Cao, 2025) only possess basic to intermediate levels of proficiency of some of these aspects. For example, few humanoids like Ameca have the potential of incorporating cognitive, sentimental or emotional states and features, such as facial expressions. In addition, social humanoids may hold humanized features and traits, such as personality as embodied by extroversion, agreeableness, conscientiousness, neuroticism and openness (Fong et al., 2003).
@@ 4. Realistic Humanoids: Other Perspectives
The comprehensive analysis of existing humanoids in Section 2, as outlined in Table 2 (Cao, 2025), and the systematic discussion on functional and nonfunctional specifications of humanoids in Section 3 disclose various gaps, limitations of existing work and the implications for building new-generation humane humanoids. Fig. 2 highlights other perspectives and directions in developing humanoid AI: humanoid humanity, mind-to-action, biomedicine, animation, virtual-real humanoid mapping and symbiosis, metaverse, humanoid-human symbiosis, generation, decentralization, standardization, and quality of services. With space limitation, this section expands Section 3 to several other perspectives and directions: cultivating humanoid humanity and subjectivity, building mind-to-action mindful and actionable humanoids, supporting transitional omnimodal humanoid modeling, and hybridizing virtual-real humanoid-human-metaverse symbiosis.
@@ 4.4. Transitional Omnimodal Humanoid Modeling
In recent years, an increasing attention has been paid to multimodal deep learning and specifically LMMs, such as visual query-answering (VQA), VLMs, VLP, image-language pretraining, image-to-text generation, and text-to-image generation. Building on Google Robotic Transformer 1 (RT-1) on multitask modeling (Brohan et al., 2023), Google Robotic Transformer 2 (RT-2) (Zitkovich et al., 2023) further supports high-capacity VLA modeling on web data and robotic vision data, and translates the learned knowledge into generalized instructions to control robotic actions. RT-2 uses the pathways language and image model (PaLI-X) and pathways language model embodied (PaLM-E) as the backbone.